# SiT: Exploring Flow and Diffusion-based Generative Models with Scalable Interpolant Transformers

### v1

@@ Appendix A Proofs
$\int_{\mathbb{R}^{d}}e^{i{\mathbf{k}}\cdot{\mathbf{x}}}\partial_{t}p_{t}({\mathbf{x}}){\mathrm{d}}{\mathbf{x}}=\int_{\mathbb{R}^{d}}{\mathbf{v}}({\mathbf{x}},t)\cdot\nabla_{{\mathbf{x}}}[e^{i{\mathbf{k}}\cdot{\mathbf{x}}}]p_{t}({\mathbf{x}}){\mathrm{d}}{\mathbf{x}}=-\int_{\mathbb{R}^{d}}\nabla_{{\mathbf{x}}}\cdot[{\mathbf{v}}({\mathbf{x}},t)p_{t}({\mathbf{x}})]e^{i{\mathbf{k}}\cdot{\mathbf{x}}}{\mathrm{d}}{\mathbf{x}}$ (23)
where $\nabla_{\mathbf{x}}\cdot[{\mathbf{v}}p_{t}]=\sum_{i=1}^{d}\frac{\partial}{\partial x_{i}}[v_{i}p_{t}]$ is the divergence operator and we take advantage of the divergence theorem and used integration by parts to get the second equality. By the properties of Fourier transform, Eq. 23 implies that $p_{t}({\mathbf{x}})$ satisfies the transport equation
@@ A.2 Proof of the SDE (4)
$d{\mathbf{X}}_{t}=[{\mathbf{v}}({\mathbf{X}}_{t},t)+\frac{1}{2}w_{t}{\mathbf{s}}({\mathbf{X}}_{t},t)]dt+\sqrt{w_{t}}d\bar{\mathbf{W}}_{t}$
@@
$\partial_{t}p_{t}({\mathbf{x}})=-\nabla_{\mathbf{x}}$ $\cdot\big{(}[{\mathbf{v}}({\mathbf{x}},t)+\frac{1}{2}w_{t}{\mathbf{s}}({\mathbf{x}},t)]p_{t}({\mathbf{x}})\big{)}+\frac{1}{2}w_{t}\Delta_{\mathbf{x}}p_{t}({\mathbf{x}})$ (25)
@@
$\implies\qquad 0$ $=-\frac{1}{2}w_{t}\nabla_{\mathbf{x}}\cdot\nabla_{\mathbf{x}}p_{t}({\mathbf{x}})+\frac{1}{2}w_{t}\Delta_{\mathbf{x}}p_{t}({\mathbf{x}})$ (27)
@@
A.3 Proof of the expression for the score in Eq. 5
@@
where we again used divergence theorem and integration by parts to get the third equality, and again the definition of the score to get the last. Comparing Eq. 33 and Eq. 37 we deduce that, when $\sigma_{t}\not=0$,
@@
$\lambda_{t}=\dot{\sigma}_{t}-\frac{\dot{\alpha}_{t}\sigma_{t}}{\alpha_{t}}$ (45)
@@
which defines the weighted score objective $\mathcal{L}_{\mathrm{s}_{\lambda}}(\theta)$. This observation is consistent with the claim made in Kingma and Gao [35] that the score objective with different monotonic weighting functions coincides with losses for different model parameterizations. In Appendix B we show that $\lambda_{t}$ corresponds to the square of the maximum likelihood weighting proposed in Song et al. [61] and Vahdat et al. [65].
Appendix B Connection with Score-based Diffusion
As shown in Song et al. [62], the reverse-time SDE from Eq. 10 is
@@ Appendix B Connection with Score-based Diffusion
we see that $2\lambda_{t}\sigma_{t}$ is precisely $\beta_{t}$, making $\lambda_{t}$ correspond to the square of maximum likelihood weighting proposed in Song et al. [62]. Further, if we plug Eq. 55 into Eq. 4 with $w_{t}=\beta_{t}$, we arrive at Eq. 51.
@@
Yet, shown in Tab. 5, we observed better performance with $\mathcal{L}_{\mathrm{s}_{\lambda}}$ for SBDM-VP, as the blowing up $\lambda_{t}$ near $t=0$ will compensate for the diminishing gradient inside the squared norm, where $\mathcal{L}_{\mathrm{v}}$ would simply experience gradient explosion resulted from $\dot{\sigma}_{t}$. The behavior is different for the Linear and GVP interpolant, where the source of instability is $\alpha^{-1}_{t}$ near $t=1$. We note $\mathcal{L}_{\mathrm{v}}$ is stable since $\alpha^{-1}_{t}$ gets cancelled out inside the squared norm, while in $\mathcal{L}_{\mathrm{s}_{\lambda}}$ it remains in $\lambda_{t}$ outside the norm.
Appendix C Sampling with Guidance
Let $p_{t}({\mathbf{x}}|{\mathbf{y}})$ be the density of ${\mathbf{x}}_{t}=\alpha_{t}{\mathbf{x}}_{*}+\sigma_{t}{\boldsymbol{\varepsilon}}$ conditioned on some extra variable ${\mathbf{y}}$. By argument similar to the one given in Sec. A.1, it is easy to see that $p_{t}({\mathbf{x}}|{\mathbf{y}})$ satisfies the transport equation (compare Eq. 24)
@@ Appendix C Sampling with Guidance
Proceeding as in Sec. A.3 and Sec. A.4, it is also easy to see that the score ${\mathbf{s}}({\mathbf{x}},t|{\mathbf{y}})=\nabla_{{\mathbf{x}}}\log p_{t}({\mathbf{x}}|{\mathbf{y}})$ is given by (compare Eq. 5)
@@
Appendix D Sampling with ODE and SDE
@@
SiT-XL 7000 9.35 / 8.61 6.38 / 6.32 126.06 / 131.65 0.67 / 0.68 0.68 / 0.67
In the main body of the paper, we used a Heun integrator for solving the ODE in Eq. 2 and an Euler-Maruyama integrator for solving the SDE in Eq. 4. We summarize all results in Tab. 1, and present the implementations below.
@@
     ${\mathbf{s}}_{\theta}\leftarrow$ convert from ${\mathbf{v}}_{\theta}$ following Sec. A.4$\triangleright$ Obtain $\nabla_{\mathbf{x}}\log p_{t}({\mathbf{x}})$ in Eq. 4
@@
It is feasible to use either a velocity model ${\mathbf{v}}_{\theta}$ or a score model ${\mathbf{s}}_{\theta}$ in applying the above two samplers. If learning the score for the deterministic Heun sampler, we could always convert the learned ${\mathbf{s}}_{\theta}$ to ${\mathbf{v}}_{\theta}$ following Sec. A.4. However, as there exists potential numerical instability (depending on interpolants) in $\dot{\sigma}_{t}$, $\alpha_{t}^{-1}$ and $\lambda_{t}$, it’s recommended to learn ${\mathbf{v}}_{\theta}$ in sampling with deterministic sampler instead of ${\mathbf{s}}_{\theta}$. For the stochastic sampler, it’s required to have both ${\mathbf{v}}_{\theta}$ and ${\mathbf{s}}_{\theta}$ in integration, so we always need to convert from one (either learning velocity or score) to obtain the other. Under this scenario, the numerical issue from Sec. A.4 can only be avoided by clipping the time interval near $t=0$. Empirically we found clipping the interval by $h=0.04$ and doing a long last step from $t=0.04$ to 00 can greatly benefit the performance. A detailed summary of sampler configuration is provided in Appendix E.
Additionally, we could replace ${\mathbf{v}}_{\theta}$ and ${\mathbf{s}}_{\theta}$ by ${\mathbf{v}}_{\theta}^{\zeta}$ and ${\mathbf{s}}_{\theta}^{\zeta}$ presented in Appendix C as inputs of the two samplers and enjoy the performance improvements coming along with guidance. As guidance requires evaluating both conditional and unconditional model output in a single step, it will impose twice the computational cost when sampling.
Comparison between DDPM and Euler-Maruyama
We primarily investigate and report the performance comparison between DDPM and Euler-Maruyama samplers. We set our Euler sampler’s number of steps to be 250 to match that of DDPM during evaluation. This comparison is made direct and fair, as the DDPM method is equivalent to a discretized Euler’s method.
Comparison between DDIM and Heun
Higher order solvers
The performances of an adaptive deterministics dopri5 solver and a second order stochastic Heun Sampler [32] are also tested. For dopri5, we set atol and rtol to 1e-6 and 1e-3, respectively; for Heun, we again maintain the NFE to be 250 to match that of DDPM. In both solvers we do not observe performance increment; under the CFG scale of $\zeta=1.5$, dopri5 and stochastic Heun gives FID-50K of 2.15 and 2.07, respectively.
We also note that our models are compatible with other samplers [36, 44] specifically tuned to diffusion models as well as sampling distillation [54, 63]. We do not include the evaluations of those methods in our work for the sake of apples-to-apples comparison with the DDPM model, and we leave the investigation of potential performance improvements to future work.
Appendix E Additional Implementation Details
We implemented our models in JAX following the DiT PyTorch codebase by Peebles and Xie [48] [^4] https://github.com/facebookresearch/DiT, and referred to Albergo et al. [3] [^5] https://github.com/malbergo/stochastic-interpolants, Song et al. [62] [^6] https://github.com/yang-song/score_sde, and Dockhorn et al. [20] [^7] https://github.com/nv-tlabs/CLD-SGM for our implementation of the Euler-Maruyama sampler. For the Heun sampler, we directly used the one from diffrax [33] [^8] https://github.com/patrick-kidger/diffrax, a JAX-based numerical differential equation solver library.

### v2

@@ Appendix 0.A Proofs
$\int_{\mathbb{R}^{d}}e^{i{\mathbf{k}}\cdot{\mathbf{x}}}\partial_{t}p_{t}({\mathbf{x}}){\mathrm{d}}{\mathbf{x}}$ $=\int_{\mathbb{R}^{d}}{\mathbf{v}}({\mathbf{x}},t)\cdot\nabla_{{\mathbf{x}}}[e^{i{\mathbf{k}}\cdot{\mathbf{x}}}]p_{t}({\mathbf{x}}){\mathrm{d}}{\mathbf{x}}$ (22)
$=-\int_{\mathbb{R}^{d}}\nabla_{{\mathbf{x}}}\cdot[{\mathbf{v}}({\mathbf{x}},t)p_{t}({\mathbf{x}})]e^{i{\mathbf{k}}\cdot{\mathbf{x}}}{\mathrm{d}}{\mathbf{x}}$ (23)
where $\nabla_{\mathbf{x}}\cdot[{\mathbf{v}}p_{t}]=\sum_{i=1}^{d}\frac{\partial}{\partial x_{i}}[v_{i}p_{t}]$ is the divergence operator and we used integration by parts to get the second equality. By the properties of Fourier transform, Eq. 23 implies that $p_{t}({\mathbf{x}})$ satisfies the transport equation
@@ 0.A.2 Proof of the SDE (4)
$d{\mathbf{X}}_{t}=[{\mathbf{v}}({\mathbf{X}}_{t},t)-\frac{1}{2}w_{t}{\mathbf{s}}({\mathbf{X}}_{t},t)]dt+\sqrt{w_{t}}d\bar{\mathbf{W}}_{t}$
@@
$\partial_{t}p_{t}({\mathbf{x}})=-\nabla_{\mathbf{x}}$ $\cdot\big{(}[{\mathbf{v}}({\mathbf{x}},t)-\frac{1}{2}w_{t}{\mathbf{s}}({\mathbf{x}},t)]p_{t}({\mathbf{x}})\big{)}-\frac{1}{2}w_{t}\Delta_{\mathbf{x}}p_{t}({\mathbf{x}})$ (25)
@@
$\implies\qquad 0$ $=\frac{1}{2}w_{t}\nabla_{\mathbf{x}}\cdot\nabla_{\mathbf{x}}p_{t}({\mathbf{x}})-\frac{1}{2}w_{t}\Delta_{\mathbf{x}}p_{t}({\mathbf{x}})$ (27)
@@
0.A.3 Proof of the expression for the score in Eq. 5
@@
where we used integration by parts to get the third equality, and again the definition of the score to get the last.
Comparing Eq. 33 and Eq. 37 we deduce that, when $\sigma_{t}\not=0$,
@@
$\lambda_{t}=\dot{\sigma}_{t}-\frac{\dot{\alpha}_{t}\sigma_{t}}{\alpha_{t}}.$ (45)
@@
which defines the weighted score objective $\mathcal{L}_{\mathrm{s}_{\lambda}}(\theta)$. This observation is consistent with the claim made in [36] that the score objective with different monotonic weighting functions coincides with losses for different model parameterizations. In Appendix 0.B we show that $\lambda_{t}$ corresponds to the square of the maximum likelihood weighting proposed in [63] and [66].
0.A.5 Proof for the optimal $w_{t}$ for tightening the KL bound
Lemma 2.22 in [2] asserts that:
$D_{\mathrm{KL}}(p({\mathbf{x}})\|p_{\theta}({\mathbf{x}}))\leq\frac{1}{2}\int_{0}^{1}w_{t}^{-1}\int_{\Omega}|b({\mathbf{x}},t)-b_{\theta}({\mathbf{x}},t)|^{2}p_{t}({\mathbf{x}}){\mathrm{d}}t{\mathrm{d}}{\mathbf{x}}$ (51)
where $p({\mathbf{x}})$ denotes the true data distribution, $p_{\theta}({\mathbf{x}})$ denotes the approximated data distribution by our model at time $t=0$, and $p_{t}$ corresponds to the marginal density in Sec. 0.A.2. We further use $b$ and $b_{\theta}$ to refer to the ground truth and approximated drift for the reverse SDE, respectively; that is, $b({\mathbf{x}},t)=v({\mathbf{x}},t)-\frac{1}{2}w_{t}s({\mathbf{x}},t)$. Following Sec. 0.A.4, $b({\mathbf{x}},t)$ can be expressed in terms of $v({\mathbf{x}},t)$
$b({\mathbf{x}},t)=(1+\frac{1}{2}\frac{w_{t}}{\lambda_{t}\sigma_{t}})v({\mathbf{x}},t)-\frac{1}{2}\frac{w_{t}\dot{\alpha}_{t}}{\alpha_{t}\lambda_{t}\sigma_{t}}x$ (52)
and similarly for $b_{\theta}({\mathbf{x}},t)$. Plug back, Eq. 51 becomes
$D_{\mathrm{KL}}(p({\mathbf{x}})\|p_{\theta}({\mathbf{x}}))\leq\frac{1}{2}\int_{0}^{1}w_{t}^{-1}(1+\frac{1}{2}\frac{w_{t}}{\lambda_{t}\sigma_{t}})^{2}\int_{\Omega}|v({\mathbf{x}},t)-v_{\theta}({\mathbf{x}},t)|^{2}p_{t}({\mathbf{x}}){\mathrm{d}}t{\mathrm{d}}{\mathbf{x}}$ (53)
Since $v({\mathbf{x}},t)=\mathbb{E}[\dot{\mathbf{x}}|{\mathbf{x}}_{t}={\mathbf{x}}]$ from Eq. 1, we have
$\int_{\Omega}|v({\mathbf{x}},t)-v_{\theta}({\mathbf{x}},t)|^{2}p_{t}({\mathbf{x}}){\mathrm{d}}{\mathbf{x}}$ $=\mathbb{E}[|v({\mathbf{x}},t)-v_{\theta}({\mathbf{x}},t)|^{2}]$
$\leq\mathbb{E}[|\dot{\mathbf{x}}-v_{\theta}({\mathbf{x}},t)|^{2}]\equiv\mathcal{L}_{t}$ (54)
where $\mathcal{L}_{t}$ is the loss at time $t$ of our model after optimization. With Eq. 55 we can further simplify Sec. 0.A.5 to be
$D_{\mathrm{KL}}(p({\mathbf{x}})\|p_{\theta}({\mathbf{x}}))\leq\frac{1}{2}\int_{0}^{1}w_{t}^{-1}(1+\frac{1}{2}\frac{w_{t}}{\lambda_{t}\sigma_{t}})^{2}\mathcal{L}_{t}{\mathrm{d}}t$ (55)
We note the minimum of the integrand in Eq. 55 is achieved at $w_{t}=2\lambda_{t}\sigma_{t}$ with a value of $2\frac{\mathcal{L}_{t}}{\lambda_{t}\sigma_{t}}$. We note that such $w_{t}$ is the exact choice of $w_{t}^{\mathrm{KL}}$ in Sec. 2.4.
For SBDM-VP interpolant, such $w_{t}^{\mathrm{KL}}$ coincides with $\beta_{t}$ in [64], and is well defined and positive everywhere on $[0,1]$. For GVP and Linear interpolant however, this diffusion coefficient is zero at $t=0$ and infinity at $t=1$. Since $\sigma_{t}=O(t)$ at $t=0$, the integrand $2\frac{\mathcal{L}_{t}}{\lambda_{t}\sigma_{t}}$ is not integrable at $t=0$, making the bound in Eq. 55 trivially $\infty$ unless $\lim_{t\to 0}\mathcal{L}_{t}=0$.
We note the bound proposed in Eq. 55 does not account for the cost of time-integration of the SDE. Assuming that the non-uniform integration step $\Delta t$ one must take to maintain a given precision is inversely proportional to $w_{t}$, that is $\Delta t=O(w_{t}^{-1})$, we can account for the integration cost by adding a term to Eq. 55
$D_{\mathrm{KL}}(p({\mathbf{x}})\|p_{\theta}({\mathbf{x}}))\leq\frac{1}{2}\int_{0}^{1}w_{t}^{-1}\left(1+\frac{1}{2}\frac{w_{t}}{\lambda_{t}\sigma_{t}}\right)^{2}\mathcal{L}_{t}{\mathrm{d}}t+\eta\int_{0}^{1}w_{t}{\mathrm{d}}t$ (56)
where $\eta>0$ is a parameter that controls the integration error: the higher the $\eta$, the smaller the cost but the higher the error, and vice-versa. The minimum of the integrand in Eq. 56 is
$\min_{w_{t}}\left(w_{t}^{-1}\left(1+\frac{1}{2}\frac{w_{t}}{\lambda_{t}\sigma_{t}}\right)^{2}\mathcal{L}_{t}+\eta w_{t}\right)=\frac{2\mathcal{L}_{t}}{\lambda_{t}\sigma_{t}}\left(1+\sqrt{\frac{4\eta\lambda_{t}^{2}\sigma_{t}^{2}+\mathcal{L}_{t}}{\mathcal{L}_{t}}}\right)$ (57)
and it is achieved at
$w_{t}=2\lambda_{t}\sigma_{t}\sqrt{\frac{\mathcal{L}_{t}}{4\eta\lambda_{t}^{2}\sigma_{t}^{2}+\mathcal{L}_{t}}}$ (58)
This is exactly the choice of $w_{t}^{\mathrm{KL},\eta}$ in Sec. 2.4. We note that such diffusion coefficient is well defined everywhere on $[0,1]$ if $\mathcal{L}_{t}$ is also well defined everywhere, and as $t\to 1$, it approaches a finite limit at $\sqrt{\frac{\mathcal{L}_{t\to 1}}{2\eta}}$.
We also note that the integrand in Eq. 56 would still be $\infty$ at time 00 given the $\frac{1}{\lambda_{t}\sigma_{t}}$, unless $\lim_{t\to 0}\mathcal{L}_{t}=0$. Theoretically, this is not an unreasonable assumption for both coefficients, as we know the closed form of $v({\mathbf{x}},0)=\mathbb{E}[\dot{\mathbf{x}}_{0}|{\mathbf{x}}_{t=0}={\mathbf{x}}]=\dot{\alpha}_{0}\mathbb{E}[{\mathbf{x}}_{*}]$ and could optimize our model $v_{\theta}({\mathbf{x}},t)$ to directly approximate this value at $t=0$. In practice, we found the numerical stability of $w_{t}^{\mathrm{KL},\eta}$ could lead to better results.
Appendix 0.B Connection with Score-based Diffusion
As shown in [64], the reverse-time SDE from Eq. 10 is
@@ Appendix 0.B Connection with Score-based Diffusion
we see that $2\lambda_{t}\sigma_{t}$ is precisely $\beta_{t}$, making $\lambda_{t}$ correspond to the square of maximum likelihood weighting proposed in [64]. Further, if we plug Eq. 63 into Eq. 4, we arrive at Eq. 59.
@@
Yet, shown in Sec. 3.1, we observed better performance with $\mathcal{L}_{\mathrm{s}_{\lambda}}$ for SBDM-VP, as the blowing up $\lambda_{t}$ near $t=0$ will compensate for the diminishing gradient inside the squared norm, where $\mathcal{L}_{\mathrm{v}}$ would simply experience gradient explosion resulted from $\dot{\sigma}_{t}$. The behavior is different for the Linear and GVP interpolant, where the source of instability is $\alpha^{-1}_{t}$ near $t=1$. We note $\mathcal{L}_{\mathrm{v}}$ is stable since $\alpha^{-1}_{t}$ gets cancelled out inside the squared norm, while in $\mathcal{L}_{\mathrm{s}_{\lambda}}$ it remains in $\lambda_{t}$ outside the norm.
Appendix 0.C Sampling with Guidance
Let $p_{t}({\mathbf{x}}|{\mathbf{y}})$ be the density of ${\mathbf{x}}_{t}=\alpha_{t}{\mathbf{x}}_{*}+\sigma_{t}{\boldsymbol{\varepsilon}}$ conditioned on some extra variable ${\mathbf{y}}$. By argument similar to the one given in Sec. 0.A.1, it is easy to see that $p_{t}({\mathbf{x}}|{\mathbf{y}})$ satisfies the transport equation (compare Eq. 24)
@@ Appendix 0.C Sampling with Guidance
Proceeding as in Sec. 0.A.3 and Sec. 0.A.4, it is also easy to see that the score ${\mathbf{s}}({\mathbf{x}},t|{\mathbf{y}})=\nabla_{{\mathbf{x}}}\log p_{t}({\mathbf{x}}|{\mathbf{y}})$ is given by (compare Eq. 5)
@@
Appendix 0.D Sampling with ODE and SDE
@@
SiT-XL 7000 9.35 / 8.26 6.38 / 6.32 126.06 / 131.65 0.67 / 0.68 0.68 / 0.67
In the main body of the paper, we used a second order Heun integrator for solving the ODE in Eq. 2 and a first order Euler-Maruyama integrator for solving the SDE in Eq. 4. We summarize all results in Tab. 8, and present the implementations below.
It is feasible to use either a velocity model ${\mathbf{v}}_{\theta}$ or a score model ${\mathbf{s}}_{\theta}$ in applying the above two samplers. If learning the score for the deterministic Heun sampler, we could always convert the learned ${\mathbf{s}}_{\theta}$ to ${\mathbf{v}}_{\theta}$ following Sec. 0.A.4. However, as there exists potential numerical instability (depending on interpolants) in $\dot{\sigma}_{t}$, $\alpha_{t}^{-1}$ and $\lambda_{t}$, it’s recommended to learn ${\mathbf{v}}_{\theta}$ in sampling with deterministic sampler instead of ${\mathbf{s}}_{\theta}$. For the stochastic sampler, it’s required to have both ${\mathbf{v}}_{\theta}$ and ${\mathbf{s}}_{\theta}$ in integration, so we always need to convert from one (either learning velocity or score) to obtain the other. Under this scenario, the numerical issue from Sec. 0.A.4 can only be avoided by clipping the time interval near $t=0$. Empirically we found clipping the interval by $h=0.04$ and doing a long last step from $t=0.04$ to 00 can greatly benefit the performance. A detailed summary of sampler configuration is provided in Appendix 0.E.
Additionally, we could replace ${\mathbf{v}}_{\theta}$ and ${\mathbf{s}}_{\theta}$ by ${\mathbf{v}}_{\theta}^{\zeta}$ and ${\mathbf{s}}_{\theta}^{\zeta}$ presented in Appendix 0.C as inputs of the two samplers and enjoy the performance improvements coming along with guidance. As guidance requires evaluating both conditional and unconditional model output in a single step, it will impose twice the computational cost when sampling.
We also note that our models are compatible with more advanced samplers [37, 46]. We do not include the evaluations of those samplers in our work for the sake of direct comparison with the DDPM model, and we leave the investigation of potential performance improvements to future work.
Comparison between DDPM and Euler-Maruyama
We primarily investigate and report the performance comparison between DDPM and Euler-Maruyama samplers. We set our Euler sampler’s number of steps to be 250 to match that of DDPM during evaluation. This comparison is made direct and fair, as the DDPM method can be viewed as a discretized version of Euler’s method.
Comparison between DDIM and Heun
@@
     ${\mathbf{s}}_{\theta}\leftarrow$ convert from ${\mathbf{v}}_{\theta}$ following Sec. 0.A.4$\triangleright$ Obtain $\nabla_{\mathbf{x}}\log p_{t}({\mathbf{x}})$ in Eq. 4
@@
Appendix 0.E Additional Implementation Details
We implemented our models in JAX following the DiT PyTorch codebase by [50] [^4] https://github.com/facebookresearch/DiT, and referred to [2] [^5] https://github.com/malbergo/stochastic-interpolants, [64] [^6] https://github.com/yang-song/score_sde, and [20] [^7] https://github.com/nv-tlabs/CLD-SGM for our implementation of the Euler-Maruyama sampler. For the Heun sampler, we directly used the one from diffrax [34] [^8] https://github.com/patrick-kidger/diffrax, a JAX-based numerical differential equation solver library.
Architectural Configurations
We follow the identical transformer architectures in DiT and have four different configurations: SiT-{S,B,L,XL}, varying in model size (parameters) and compute (flops). A detailed summarization is presented below.
Model Layers $N$ Hidden size $d$ Heads
SiT-S 12 384 6
SiT-B 12 768 12
SiT-L 24 1024 16
SiT-XL 28 1152 16